Suppose that the moment generating function of X is M(t) 1-2t . Find E[X]rounded to nearest .xx.
Answer #1

solution: Ceiven mit off is Mit)= 1-t St E(X) = m (O) e It derivative of mit & put to - (m(es) Eco = 2 = (1-25) to 2 = [1-2(0